# Diffusion Actor-Critic with Entropy Regulator

### v2

@@ 1 Introduction
In summary, the key contributions of this paper are the following: 1) We propose to consider the reverse process of the diffusion model as a novel policy function. The objective function of the diffusion policy is to maximize the expected Q-value and thus achieve policy improvement. 2) We propose a method for estimating the entropy of diffusion policy. The estimated value is utilized to achieve an adaptive adjustment of the exploration level of the diffusion policy, thus improving the policy performance. 3) We evaluate the efficiency and generality of our method on the popular MuJoCo benchmarking. Compared with DDPG silver2014deterministic , TD3 fujimoto2018addressing , PPO schulman2017proximal , SAC haarnoja2018soft , DSAC duan2023dsac ; duan2021distributional , and TRPO schulman2015trust , our approach achieves the SOTA performance. In addition, we demonstrate the superior representational capacity of our algorithm through a specific multi-goal task. 4) We provide the DACER code written in PyTorch and JAX in the attachment file to facilitate future researchers to follow our work.
@@ 2 Related Work
Online RL, characterized by real-time environment interaction, contrasts with offline RL’s dependence on pre-existing datasets. To date, only two studies have delved into integrating online RL with diffusion models. Yang et al. yang2023policy pioneered this approach by using the action gradient method for action updates in replay buffers, facilitating diffusion policy improvement and theoretically proving its convergence. However, this method diverges from traditional policy evaluation and policy improvement in RL, incorporating changes to the core of RL that hinder compatibility with established algorithms and is also time-inefficient. Psenka et al. psenka2023learning introduced Q-score matching (QSM) for off-policy online RL, moving away from conventional behavioral cloning by leveraging the relationship between a score of policy and the action gradient Q-function in diffusion models. The QSM approach is also difficult to combine with existing RL algorithms. Because, in contrast with the traditional RL policy network that maximizes the expected reward, its objective function aims to minimize the discrepancy between the derivative of the Q function concerning action and the derivative of the log strategy probability concerning action. Moreover, the multimodal aspects of the diffusion policy remain underexplored. The performance of the above two methods in MuJoCo is lower than that of SAC, so we do not compare with them in experiments.
@@ 3.1 Online Reinforcement Learning
$J_{\pi}=\mathbb{E}_{(s_{i\geq t},a_{i\geq t})\sim\pi}\Big{[}\sum_{i=t}^{\infty}\gamma^{i-t}r_{i}\Big{]},$ (1)
@@
$Q(s,a)=\mathbb{E}_{\pi}\left[\sum_{i=t}^{\infty}\gamma^{i-t}r_{i}\right].$ (2)
@@
In practical applications, neural networks are often used to parameterize both the policy and value functions, represented by $\pi_{\theta}$ and $Q_{\phi}$, respectively. These functions are refined through the application of gradient descent methods aimed at reducing the loss functions for both the actor, $\mathcal{L}_{\pi}(\theta)$, and the critic, $\mathcal{L}_{q}(\phi)$. These loss functions are structured based on the principles outlined in (4) and (3).
@@ 4.3 Diffusion Policy with Entropy
Iterative optimization continues until parameter convergence. Based on our experimental experience in the MuJoCo environments, a general setting of $K=4$ provides a better fit to the action distribution.
According to Eq. (12), we can estimate the entropy of the action distribution corresponding to the state by
$\mathcal{H}_{\boldsymbol{s}}=-\sum_{k=1}^{K}w_{k}\log w_{k}+\sum_{k=1}^{K}w_{k}\cdot\frac{1}{2}\log\left((2\pi e)^{d}|\boldsymbol{\Sigma}_{k}|\right),$ (15)
@@ 5 Experiments
To ensure a fair comparison, we incorporated the diffusion policy as a policy approximation function within GOPS and implemented DACER with JAX, which improves training speed by 4-5 times compared to PyTorch while maintaining consistent performance. All algorithms and tasks use the same three-layer MLP neural network with GeLU hendrycks2016gaussian or Mish misra2019mish activation functions, the latter used only for the noise prediction network in the diffusion policy. Initially, we encode timestep $t$ into 16 dimensions using sinusoidal embedding vaswani2017attention , then merge this encoded result with the state $\boldsymbol{s}$ and action $\boldsymbol{a}_{t}$ during the current denoising step, and input it into the prediction noise network to generate the output. The impact of the reverse diffusion step size, $T$, on the experimental results will be examined in the ablation experiments. $T$ is set to 20 for all experiments eventually. The Adam kingma2014adam optimization method is employed for all parameter updates. In this paper, the total training step size for all experiments is set at 1.5 million, with the results of all experiments averaged over five random seeds. The CPU used for the experiment is Xeon(R) Platinum 8352V, and the GPU is NVIDIA GeForce RTX 4090. Taking Humanoid-v3 as an example, the time taken to train 1.5 million in the JAX framework is 7 hours. More detailed hyperparameters are provided in Appendix A.2 due to space limits.
@@ 5.2 Policy Representation Experiment
Figure 3: Trainining curves.
@@ 5.3 Ablation Study
In section 4.2, we propose a method using the reverse diffusion process as a policy approximator, which can be combined with the non-maximizing entropy RL algorithm. However, the diffusion policy trained without entropy exhibits poor exploratory properties, leading to suboptimal performance. Using Walker2d-v3 as an example, we compared the training curves of this method with the DACER algorithm, as shown in Fig. 3.
@@
In order to verify that using the estimated entropy to adaptively adjust the noise factor plays an important role in the final performance, we conducted the following two experiments in the Walker2d-v3 task: 1) Fixed noise factor to 0.1; 2) The noise factor starts from 0.27 and linearly decreases to 0.1 during the training process. These two values were chosen because the starting and ending noise factor for adaptive tuning in this setting is about in this range. As shown in Fig. 4, our method of adaptively adjusting the noise factor based on the estimated entropy achieves the best performance.
Figure 5: Trainining curves.
@@
We further examined the performance of the diffusion policy as the number of diffusion timesteps $T$ varied. We used the Walker2d-v3 task to plot training curves for $T=10,20,$ and $30$, as shown in Fig. 5. Experimental results indicate that a larger number of diffusion steps does not necessarily lead to better performance. Excessive diffusion steps can cause gradient explosion, significantly reducing the performance of diffusion policy. After balancing performance and computational efficiency, we selected 20 diffusion steps for all experiments.
@@ A.2 Training Details on MuJoCo tasks
Buffer warm-up size (Humanoid) 100000
@@
Target network soft-update rate $\tau$ 0.005
@@
Number of Gaussian distributions for mixing 4
Number of action samples in entropy estimation 500
@@
Noise scale $\lambda$ (Humanoid) 0.15
@@
Target entropy $-{\rm{dim}}(\mathcal{A})$

### v4

@@ 1 Introduction
In summary, the key contributions of this paper are the following: 1) We propose to consider the reverse process of the diffusion model as a novel policy function. The objective function of the diffusion policy is to maximize the expected Q-value and thus achieve policy improvement. 2) We propose a method for estimating the entropy of diffusion policy. The estimated value is utilized to achieve an adaptive adjustment of the exploration level of the diffusion policy, thus improving the policy performance. 3) We evaluate the efficiency and generality of our method on the popular MuJoCo benchmarking. Compared with DDPG silver2014deterministic , TD3 fujimoto2018addressing , PPO schulman2017proximal , SAC haarnoja2018soft , DSAC duan2023dsac ; duan2021distributional , and TRPO schulman2015trust , our approach achieves the SOTA performance. In addition, we demonstrate the superior representational capacity of our algorithm through a specific multi-goal task. 4) We provide the DACER code written in PyTorch and JAX to facilitate future researchers to follow our work.
@@ 2 Related Work
Online RL, characterized by real-time environment interaction, contrasts with offline RL’s dependence on pre-existing datasets. To date, only two studies have delved into integrating online RL with diffusion models. Yang et al. yang2023policy pioneered this approach by proposing action gradient method. This approach achieves policy improvement by updating the action in the replay buffer through the $\nabla_{a}Q$, followed by mimicry learning of the action post-update using a diffusion model. However, action gradient increased the additional training time. Furthermore, it is difficult to fully learn both the action gradient and imitation learning steps simultaneously, which also resulted in suboptimal performance of this method in MuJoCo tasks. Psenka et al. psenka2023learning proposes Q-score matching (QSM), a new methodology for off-policy reinforcement learning that leverages the score-based structure of diffusion model policies to align with the $\nabla_{a}Q$. This approach aims to overcome the limitations of simple behavior cloning in actor-critic settings by integrating the policy’s score with the Q-function’s action gradient. However, QSM needs to accurately learn $\nabla_{a}Q$ in most of the action space to achieve optimal guidance. This is difficult to accomplish, resulting in suboptimal performance of QSM.
@@ 3.1 Online Reinforcement Learning
$J_{\pi}=\mathbb{E}_{(s_{i\geq t},a_{i\geq t})\sim\pi}\Big{[}\sum_{i=t}^{\infty}\gamma^{i-t}r(s_{i},a_{i})\Big{]},$ (1)
@@
$Q(s,a)=\mathbb{E}_{\pi}\Big{[}\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})|s_{0}=s,a_{0}=a\Big{]}.$ (2)
@@
In practical applications, neural networks are often used to parameterize both the policy and value functions, represented by $\pi_{\theta}$ and $Q_{\phi}$, respectively. These functions are refined through the application of gradient descent methods aimed at reducing the loss functions for both the critic, $\mathcal{L}_{q}(\theta)=\mathbb{E}_{({s},{a},{s}^{\prime})\sim\mathcal{B}}\left[\left(r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p,a^{\prime}\sim\pi}[Q^{\pi}(s^{\prime},a^{\prime})]-Q^{\pi}_{\phi}(s,a)\right)^{2}\right]$, and the actor, $\mathcal{L}_{\pi}(\phi)=-\mathbb{E}_{s\sim d_{\pi},a\sim\pi}[Q^{\pi_{\mathrm{old}}}(s,a)]$. These loss functions are structured based on the principles outlined in (3) and (4).
@@ 4.3 Diffusion Policy with Entropy
Iterative optimization continues until parameter convergence. Based on our experimental experience in the MuJoCo environments, a general setting of $K=3$ provides a better fit to the action distribution.
According to Eq. (12), we can estimate the entropy of the action distribution corresponding to the state by huber2008entropy
$\mathcal{H}_{\boldsymbol{s}}\approx-\sum_{k=1}^{K}w_{k}\log w_{k}+\sum_{k=1}^{K}w_{k}\cdot\frac{1}{2}\log\left((2\pi e)^{d}|\boldsymbol{\Sigma}_{k}|\right),$ (15)
@@ 5 Experiments
To ensure a fair comparison, we incorporated the diffusion policy as a policy approximation function within GOPS and implemented DACER with JAX, which improves training speed by 4-5 times compared to PyTorch while maintaining consistent performance. All algorithms and tasks use the same three-layer MLP neural network with GeLU hendrycks2016gaussian or Mish misra2019mish activation functions, the latter used only for the noise prediction network in the diffusion policy. Initially, we encode timestep $t$ into 16 dimensions using sinusoidal embedding vaswani2017attention , then merge this encoded result with the state $\boldsymbol{s}$ and action $\boldsymbol{a}_{t}$ during the current denoising step, and input it into the prediction noise network to generate the output. The impact of the reverse diffusion step size, $T$, on the experimental results will be examined in the ablation experiments. $T$ is set to 20 for all experiments eventually. The Adam kingma2014adam optimization method is employed for all parameter updates. In this paper, the total training step size for all experiments is set at 1.5 million, with the results of all experiments averaged over five random seeds. The CPU used for the experiment is the AMD Ryzen Threadripper 3960X 24-Core Processor, and the GPU is NVIDIA GeForce RTX 3090Ti. Taking Humanoid-v3 as an example, the time taken to train 1.5 million in the JAX framework is 5 hours. More detailed hyperparameters are provided in Appendix A.2 due to space limits.
@@ 5.2 Policy Representation Experiment
To demonstrate the powerful multimodality of DACER, we select five points requiring multimodal policies: (0.5, 0.5), (0.5, -0.5), (-0.5, -0.5), (-0.5, 0.5), and (0, 0). For each point, we sampled 100 trajectories. The trajectories are plotted in Fig. 3. The results show that compared with DSAC, DACER exhibits strong multimodality. This also explains why only the Q-function of DACER can learn the nearly perfectly symmetrical four peaks.
Figure 3: Multi-goal multimodal experiments. We selected 5 points that require multimodal policies: (0, 0), (-0.5, 0.5), (0.5, 0.5), (0.5, -0.5), (-0.5, -0.5), and sampled 100 trajectories for each point. The top row shows the experimental results of DACER, another shows the experimental results of DSAC.
@@ 5.3 Ablation Study
In section 4.2, we propose a method using the reverse diffusion process as a policy approximator, which can be combined with the non-maximizing entropy RL algorithm. However, the diffusion policy trained without entropy exhibits poor exploratory properties, leading to suboptimal performance. Using Walker2d-v3 as an example, we compared the training curves of this method with the DACER algorithm, as shown in Fig. 4(a).
@@
In order to verify that using the estimated entropy to adaptively adjust the noise factor plays an important role in the final performance, we conducted the following two experiments in the Walker2d-v3 task: 1) Fixed noise factor to 0.1; 2) The noise factor starts from 0.27 and linearly decreases to 0.1 during the training process. These two values were chosen because the starting and ending noise factor for adaptive tuning in this setting is about in this range. As shown in Fig. 4(b), our method of adaptively adjusting the noise factor based on the estimated entropy achieves the best performance.
@@
We further examined the performance of the diffusion policy as the number of diffusion timesteps $T$ varied. We used the Walker2d-v3 task to plot training curves for $T=10,20,$ and $30$, as shown in Fig. 4(c). Experimental results indicate that a larger number of diffusion steps does not necessarily lead to better performance. Excessive diffusion steps can cause gradient explosion, significantly reducing the performance of diffusion policy. After balancing performance and computational efficiency, we selected 20 diffusion steps for all experiments.
(a) Ablation for the entropy regulator mechanism.
(b) Ablation for the noise factor modulation mechanism.
(c) Ablation for the different diffusion steps.
Figure 4: Ablation experiment curves. (a) DAC stands for not using the entropy regulator. DACER’s performance on Walker2d-v3 is far better than DAC. (b) Adaptive tuning of the noise factor based on the estimated entropy achieved the best performance compared to fixing the noise factor or using the adaptive tuning method with initial, end values followed by a linear decay method. (c) he best performance was achieved with diffusion steps equal to 20, in addition to the instability of the training process when equal to 30.
@@ 6 Conclusion
7 Acknowledgements
This study is supported by National Key R&D Program of China with 2022YFB2502901, and Tsinghua University Initiative Scientific Research Program.
@@ A.2 Training Details on MuJoCo tasks
Buffer warm-up size (Humanoid, HalfCheetah) 200000
@@
Target network soft-update rate $\rho$ 0.005
@@
Number of Gaussian distributions for mixing 3
Number of action samples in entropy estimation 200
@@
Noise scale $\lambda$ (Humanoid, HalfCheetah) 0.15
@@
Target entropy $-0.9\cdot{\rm{dim}}(\mathcal{A})$